4) Sir Paul Incorporated is a company engaged in the selling of household floor tiles. The company produced 15000 pcs of tiles on its first month of business and sells it at P50 per piece since it is the accepted market price. On its next month, the price of the materials used to create the tiles however increased and as a result, the company was forced to increase the price of the tiles it sells by 50%. The company invested on new machines that enable the workers to produce tiles faster. Production on the part of the company tripled compared to the previous month. Compute for supply elasticity. Is it elastic, inelastic or unitary? Write it beside your supply elasticity coefficient.
1
Expert's answer
2016-07-28T09:03:02-0400
Q1 = 15000 pcs of tiles, P1 = 50 per piece, Q2 = 45000 pcs, P2 = 75 per piece.
The supply elasticity coefficient is:
Es = (45000 - 15000)/(75 - 50)*(75 + 50)/(45000 + 15000) = 30000/25*125/60000 = 2.5, so the supply is elastic.